We have now 400G instead of 800G between DC2/DC3 Core backbone and DC5 Aggregators
ADVA transmission amplifiers are fine but simply no signal. No impact until DC3 <> DC5 span is not affected
Opened ticket to Sipartech supplier / provider of the dark fiber.
30/01/20 1050Z (1150LT)
Our provider has moved the cable for works and since then, we identified the cut.
We are working with them closely to fix it as soon as possible.
The technician has been dispatched and should be on-site around 3.00pm
30/01/20 1950Z (2050LT)
Technicians have fixed the problem but we still have an increased attenuation between TH2 and DC5. Technicians are still searching where it comes from.
Traffic is flowing through the redundant fibers' pair.